To recharge a 18.0 V battery, a battery charger must move 2.56x105 C of charge from the negative terminal to the positive terminal How much work is done by the battery charger? Submit Answer Tries 0/12
Answer #1

here,

the potential difference of battery , V = 18 V

charge , Q = 2.56 * 10^5 C

the work done by the battery charger , W = 0.5 * Q^2 /V

W = 0.5 * (2.56 * 10^5)^2 /18 J

W = 1.82 * 10^9 J

the work done by the battery charger is 1.82 * 10^9 J